OpenStudy (anonymous):

The largest building in the world by volume is the Boeing 747 plant in Everett, Washington. It measures approximately 634 long, 710 yards wide, and 112 high. What is its volume in cubic feet?

OpenStudy (anonymous):

Well i'm assuming that this is a huge rectangular shaped building. Do you know how to find the volume of a rectangle?

OpenStudy (anonymous):

yes its more so the conversions i'm having problems with I got 490717920 cubic feet

OpenStudy (anonymous):

Ah yeah, conversions can be tricky. Are all sides in yards?

OpenStudy (whpalmer4):

Best way to do them is as you go along: V = l * w * h \[V = (634 yards)(3 feet / 1 yard) * (710 yards)(3 feet / 1 yard) * (112 yards)(3 feet / 1 yard)\] Make sure the units cancel properly! That's a valuable check that you've set up the problem and the conversions correctly.

OpenStudy (anonymous):

sorry the units didnt come over it. its 627 meters long, 710 yards wide, 112 ft high...looking for cubic feet volume

OpenStudy (whpalmer4):

112 feet high or 112 yards?

OpenStudy (whpalmer4):

I've been there, but don't remember how tall the building was, only how expansive.

OpenStudy (anonymous):

feet. someone else posted a similar question

OpenStudy (whpalmer4):

Ah, yes, feet it is, and the building is apparently not quite the same dimensions or shape, because Boeing's figure for the volume is a bit different. So that would be: \[V = (634 yards)(3 feet / 1 yard) * (710 yards)(3 feet / 1 yard) * (112 feet)\]

OpenStudy (anonymous):

So when ever you want to conver the units, use dimensional analysis. 1 Meter = 3.281 ft \[627 Meters * \frac{ 3.281 Feet }{ 1 Meter }\] We then just need to do this for the rest of them. I don't remember the conversions, and just need to look them up each time.

OpenStudy (whpalmer4):

Notice that if you had mistakenly applied the conversion factor (written with units, as I've done) to the height, despite measuring in yards instead of feet, the units would not cancel properly: \[yards * \frac{feet}{yards} * yards * \frac{feet}{yards} * feet * \frac{feet}{yards} = \frac{feet^4}{yards} \] which is clearly not a valid unit!

OpenStudy (whpalmer4):

er, make that "despite measuring in feet instead of yards"...

OpenStudy (anonymous):

4.91* 10^8 ft ^3 was the answer I originally got that using the same conversions as yall but forgot to do scientific notation. If I wanted to change it to cubic meters would I just divide by .0283?

OpenStudy (anonymous):

sorry 4.91*10^8 is correct according to masterphyiscs

OpenStudy (whpalmer4):

Hmm, that's not the number I get...634 * 3 * 710 * 3 * 112 = 453,741,120 cubic feet Boeing's figure is 472,370,319 cubic feet.

OpenStudy (whpalmer4):

Oh, wait, you changed 634 to 627 when you restated, so now the answer is 448,731,360 cubic feet.